OpenStudy (anonymous):

Evaluate: log(x)^1/2=log1-3log2

OpenStudy (anonymous):

\[log(1)=0\] \[3 log(2)=log(2^3)=log(8)\] so \[x^\frac{1}{2}=8\]

OpenStudy (anonymous):

making x = 64

OpenStudy (anonymous):

okay thanks :)

OpenStudy (anonymous):

yw